OpenStudy (nali):

57 plz help... find the vertex, focus, and directrix of the parabola and sketch its graph x^2 + 4x + 6y - 2= 0

OpenStudy (nali):

u do not have to graph it

OpenStudy (anonymous):

a good explanation is here http://www.purplemath.com/modules/parabola.htm

OpenStudy (nali):

i already know all these things but i do not know how to apply it to this problem

OpenStudy (anonymous):

answer is here http://www.wolframalpha.com/input/?i=focus+x^2+%2B+4x+%2B+6y+-+2%3D+0 but i think to get there we have to rewrite this thing

OpenStudy (nali):

can u plz show me how they got the focus,vertex and directrix instead of just having the answers

OpenStudy (anonymous):

ok lets see if we can write it in the form \[4p(y-k)^2=(x-h)^2\]

OpenStudy (anonymous):

then we can read off the answer. so the next part is just algebra

OpenStudy (anonymous):

\[x^2 + 4x + 6y - 2= 0\] \[6y-2=-x^2-4x\] now complete the square on the right to get \[6y-2=-(x^2+4x)\] \[6y-2=-(x+2)^2+4\]

OpenStudy (anonymous):

does that part make sense?

OpenStudy (nali):

the equation is actually (x-h)^2 = 4p (y-k)

OpenStudy (anonymous):

equality is reflexive, so we are free to write it either way

OpenStudy (anonymous):

oh damn i see i made a typo!

OpenStudy (nali):

that is what is meant

OpenStudy (anonymous):

sorry about that it should be \[4p(y-k)=(x-h)^2\]

OpenStudy (anonymous):

ok so far we have \[6y-2=-(x+2)^2+4\] is that part ok? only two more steps

OpenStudy (anonymous):

\[6y-6=-(x+2)^2\] \[6(y-1)=-(x+2)^2\] \[-6(y-1)=(x+2)^2\] now we can read off the answer

OpenStudy (nali):

where did u get the +4 at the end from?

OpenStudy (anonymous):

ok lets go over that slowly focusing only on the right hand side, we had \[-x^2-4x\]

OpenStudy (anonymous):

factor out the \(-1\) get \(-(x^2+4x)\) and now we complete the square

OpenStudy (anonymous):

half of 4 is 2, so we write \(-(x-2)^2+4\) and the reason for the \(+4\) is that \[-(x+2)^2=-x^2-4x-4\] so we have actually subtracted 4 and so now we must add it back, otherwise it is no longer what we started with

OpenStudy (nali):

got it ok so what is the next step?

OpenStudy (anonymous):

for that matter we could have written \[6y-2=-(x^2+4x)\] and then \[6y-2-4=-(x+2)^2\] because if we subtract 4 from the right, we have to subtract 4 form the left either way we end up with \[6y-6=-(x+2)^2\]

OpenStudy (nali):

that is actually either but i understand it both ways

OpenStudy (nali):

actually easier*

OpenStudy (anonymous):

the we get to the form we want, namely \[4p(y-k)=(x-h)^2\] by multiplying both sides by \(-1\) to get \[-6(y-1)^2=(x+2)^2\]

OpenStudy (anonymous):

yeah probably easier since we have an equation. the important thing to remember is that because of the minus sign out front you are subtracting 4, not adding it

OpenStudy (anonymous):

now we should be able to read the answer from the form we have we know it is a parabola that opens down. we knew that from the start

OpenStudy (anonymous):

vertex is \((h,k)\) in this case it is \((-2,1)\)

OpenStudy (nali):

is that another typo when u say -6 (y-1) ^2 above

OpenStudy (anonymous):

yeah something is the matter with me

OpenStudy (anonymous):

this is the form we have and it is what we want \[-6(y-1)=(x+2)^2\]

OpenStudy (anonymous):

so we can read off the vertex right?

OpenStudy (anonymous):

then we know that it is a parabola opening down with vertex at \((-2,1)\) now we can find the focus

OpenStudy (anonymous):

\[4p=-6\implies p=-\frac{3}{2}\] we go down \(-\frac{3}{2}\) units from \((-2,1)\) and end up at \((-2,-\frac{1}{2})\) so that is the focus

OpenStudy (nali):

thanks so much i got it from this point on

OpenStudy (anonymous):

and finally the directrix is \(\frac{3}{2}\) units above \((-2,1)\) namely \(y=1+\frac{3}{2}=\frac{5}{2}\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

and sorry about the typos

OpenStudy (nali):

it is ok